# How Transformers Solve Propositional Logic Problems: A Mechanistic Analysis

### v2

@@ 4.3 Circuit analysis
In this section, we discuss properties of the reasoning circuit of Mistral-7B. The order by which we present the results will be from the coarser- to the finer-grained, roughly following the process which we discovered the circuit; we believe this adds greater transparency to the circuit discovery process. We delay the more involved (and complete) set of experimental results to Appendix C.
4.3.1 QUERY-based patching: discovering the important attention heads
@@
We initiate our analysis with QUERY patching, following the same procedure as detailed in sub-section 4.2. In this set of experiments, we discover the main attention heads responsible for processing the context and performing inference as introduced in the beginning of this Section.
(a) Residual stream patching
(b) Attention block patching
(c) MLP patching
Figure 4: Query patching at the level of residual streams, attention blocks and MLPs. Highly localized processing of QUERY is observed: a sharp transition occurs at layer 13 in (a), and in (b), only a sparse set of attention blocks play a major role in this subtask. Furthermore, (c) shows that the MLPs play a limited role in this subtask (besides MLP0). (Please zoom in for the details)
Why is QUERY-based patching important to reasoning circuit discovery? To answer this question, there are two points to emphasize first. (1) We know that to solve the reasoning problem, the QUERY token is critical to initiating the reasoning chain: without it, the rules and facts are completely useless; with it, the reasoner can then proceed to identify the relevant rules and facts to predict the answer. (2) The prompt pairs differ only by the QUERY token. Based on (1) and (2), we know that if performing the aforementioned QUERY-based causal intervention on a model component leads to a large intervened logit difference (i.e. it alters the model’s “belief”), then this component must be integral to the reasoning circuit, because the component is now identified to be QUERY-sensitive and has causal influence on (parts of) the model’s reasoning actions.
High-level interventions. We begin by presenting higher-level patching results in Figure 4, where we intervene at the level of residual streams, attention blocks, and MLPs. We can draw a few insights from these results:
  1. 1.
    A sharp transition of “QUERY processing” occurs from layer 12 to layer 13 (indexed from 0) in Figure 4(a) and (b).
  2. 2.
    Figure 4(b) shows that a small set of attention blocks are observed to be significant for the “belief altering” action, namely those in layers 9, 12, 13, 14, 16 and 19.
  3. 3.
@@
    The MLPs, shown in Figure 4(c), play little role in this circuit, except for MLP-0. However, MLP-0 had been observed to act more as a “nonlinear token embedding” than a complex high-level processing unit (Wang et al., 2023). In the rest of this section, we primarily devote our analysis to the attention heads, and leave the exact role of the MLPs to future work.
@@
Remark. In Figure 4 and the rest, we adopt a calibrated version of the logit difference (see Appendix C): the closer to 1, the more significant the component is in “altering” the “belief” of the model on the selected subtask.
Attention-head interventions. Figure 4 helps us locate a small set of attention blocks which are important to the task. However, these results alone still leave us with far too many components to examine in detail. Therefore, we run a set of finer-grained experiments, intervening on the attention heads (over the relevant context). The results are shown in Figure 5. We find that, interestingly, only a very small set of attention heads are central to the “belief altering” of the LLM. More specifically, in Figure 5(a), only attention heads (12,9), (13,11), (14,24;26), (16,0;12;14), (17,25), (19,8;9;16) are observed with relatively high intervened logit differences.
Figure 5: Attention head patching, highlighting the ones with the highest intervened logit difference; $x$-axis is the head index. (a) shows single-head patching, and (b) shows a coarser-grained head patching in groups. In (b), we only highlight the head groups that are not captured well by (a).
We note that Grouped-Query Attention used by Mistral-7B adds subtlety to the analysis [^3] In Mistral-7B-v0.1, each attention layer has 8 key and value activations, and 32 query activations. Therefore, heads $(\ell,h\times 4)$ to $(\ell,h\times 4+3)$ share the same key and value activation.: patching a single head might not yield a high logit difference since other heads in the same group (which possibly perform a similar function) could overwhelm the patched head and maintain the model’s previous “belief”. Therefore, we also run a coarser-grained experiment which simultaneously patches the attention heads sharing the same key and value activations, shown in Figure 5(b). This experiment reveals that heads belonging to the group (9, 24 - 27) also have high intervened logit difference.
4.3.2 Finer examinations of the attention heads
@@
Attention-head sub-component patching (QUERY-based patching). We now aim to understand why the attention heads identified in the last sub-section are important. For now, we continue with QUERY altering in the prompt pairs. Through intervening on the sub-components of each attention head, namely their value, key, and query, and through examining details of their attention weights, we find that there are roughly four types of attention heads. We show the results in Figure 6:
  1. 1.
@@
    Queried-rule locating head. Attention head (12,9)’s query activation has a large intervened logit difference according to Figure 6(a), therefore, its query and attention patterns are QUERY-dependent and contribute to altering the model’s “belief”. Furthermore, at the QUERY position, we find that on average, its attention weight is above 90% at the “conclusion” variable of the rule being queried. In other words, it is responsible for locating the queried rule, and storing that rule’s information at the QUERY position.444(9,25;26), (14,24;26) exhibit similar tendencies, albeit with smaller intervened logit differences.
  2. 2.
@@
    Queried-rule mover head. Attention head (13,11)’s value activations have large intervened logit difference, and intriguingly, its query and key activations do not share that tendency. This already suggests that its attention pattern performs a fixed action on both the original and altered prompts, and only the value information is sensitive to QUERY. Furthermore, within the relevant context (excluding the 6 in-context examples given), head (13,11) assigns above 50% attention weight to the QUERY position, and its attention weight at QUERY is about 10 times larger than the second largest one on average. Recalling the role of layer 12, we find evidence that head (13,11) moves the QUERY and queried-rule information to the “:” position.555(15,8), (16,0) also appear to belong to this type, albeit with smaller intervened logit difference.
  3. 3.
@@
    Fact processing heads. Attention heads (16,12;14)’s query activations have large intervened logit differences. Within the relevant context, they place greater than 56%, and 70% of their attention respectively in the fact section of the context (starting from “Fact” and ending on “.” before “Question”).
  4. 4.
@@
    Decision head. Attention head (19,8)’s query activations have large intervened logit differences. Its attention pattern suggests that it is a “decision” head: within the relevant context, when the model is correct, the head’s top-2 attention weights are always on the correct starting node of the queried rule and the correct variable in the fact section, and the two token positions occupy more than 60% of its total attention in the relevant context on average. In other words, it already has the answer.666(17,25), (19,9;16) exhibit similar tendencies, albeit with smaller intervened logit differences.
@@
Due to space limits, we delay our more involved experiments to Appendix C, including closer inspection of the attention patterns of the attention head families (App. C.2.2), and further insights and validating evidence for the queried-rule locating heads (App. C.3) and the fact-processing and decision heads (App. C.4).
@@ 5 Conclusion
In future work, we plan to expand our analysis to uncover possibly greater families of model components which are important to an LLM’s reasoning circuit on simple logic problems resembling the one in this paper. We also plan to investigate if similar circuits are employed when an LLM solves GSM-like problems. Finally, we are interested in understanding the circuit in architecturally different LLMs such Mistral 8x7B, which uses mixture of experts.
@@ B.2.1 Transformer definition
The architecture definition follows that of GPT-2 closely.
@@
Define input ${\bm{x}}=(x_{1},x_{2},...,x_{t})\in\mathbb{N}^{t}$, a sequence of tokens with length $t$. It is converted into a sequence of (trainable) token embeddings ${\bm{X}}_{token}=({\bm{e}}(x_{1}),{\bm{e}}(x_{2}),...,{\bm{e}}(x_{t}))\in\mathbb{R}^{d_{in}\times t}$. Adding to it the (trainable) positional embeddings ${\bm{P}}=({\bm{p}}_{1},{\bm{p}}_{2},...,{\bm{p}}_{t})\in\mathbb{R}^{d_{in}\times t}$, we form the zero-th layer embedding of the transformer ${\bm{X}}_{0}=({\bm{e}}(x_{1})+{\bm{p}}_{1},...,{\bm{e}}(x_{t})+{\bm{p}}_{t})$. The input is processed by the attention blocks as follows.
Let the model have $L$ layers and $H$ heads. For layer index $\ell\in[L]$ and head index $j\in[H]$, attention head $\mathcal{A}_{\ell,j}$ is computed by $\mathcal{A}_{\ell,j}({\bm{X}}_{\ell-1})=\mathcal{S}\left(\text{causal}\left[\tilde{{\bm{X}}}_{\ell-1}{\bm{Q}}_{\ell,j}^{T}{\bm{K}}_{\ell,j}\tilde{{\bm{X}}}_{\ell-1}^{T}\right]/\sqrt{d_{H}}\right)\tilde{{\bm{X}}}_{\ell-1}{\bm{V}}_{\ell}^{T}$, with $\tilde{{\bm{X}}}_{\ell-1}=\text{LayerNorm}({\bm{X}}_{\ell-1})$, $\mathcal{S}(\cdot)$ being the softmax operator, $\text{causal}[\cdot]$ the causal mask operator. The output of the attention block is ${\bm{X}}_{\ell}={\bm{X}}_{\ell-1}+\text{Concat}[\mathcal{A}_{\ell,1}({\bm{X}}_{\ell-1}),...,\mathcal{A}_{\ell,H}({\bm{X}}_{\ell-1})]{\bm{W}}_{O,\ell}^{T}$, with ${\bm{W}}_{O,\ell}$ the square output matrix (with bias). Finally, we apply an affine classifier (with softmax) ${\bm{f}}({\bm{x}})=\mathcal{S}(\tilde{{\bm{X}}}_{L,t}{\bm{W}}_{class}^{T}+{\bm{b}}_{class})$ to predict the next word.
In this paper, we set the hidden space embedding to 768.

### v3

@@ 4.3 Circuit analysis
In this section, we discuss properties of the reasoning circuit of Mistral-7B. The order by which we present the results will roughly follow the process which we discovered and verified the circuit; we believe this adds greater transparency to the circuit analysis process. We delay the more involved (and complete) set of experimental results to Appendix C.
4.3.1 Circuit discovery: necessity-based search for model components
@@
We initiate our analysis with QUERY-based patching, following the same procedure as detailed in §4.2. In this set of experiments, we discover the main attention heads responsible for processing the context and performing inference as introduced in the beginning of this Section.
@@
High-level interventions. Figure 4(a) helps us locate a small set of attention heads which are central to the “belief altering” of the LLM. More specifically, only attention heads (12,9), (13,11;22), (14,24;26), (16,0;12;14), (17,25), (19,8;9;16), and (9,25;26) [^4] We discover these two heads with a slightly different patching experiment, presented in Appendix C.3.2. are observed with relatively high intervened logit differences. As for the MLPs, shown in Figure 4(b), play little role in this circuit, except for MLP-0. However, MLP-0 had been observed to act more as a “nonlinear token embedding” than a complex high-level processing unit (Wang et al., 2023). In the rest of this section, we primarily devote our analysis to the attention heads, and leave the exact role of the MLPs to future work.
Figure 4: Attention head and MLP output patching results (over all token positions in relevant context) in (a) and (b) respectively. The closer to 1, the greater the causal influence a component has on the LLM’s correct inference. We highlight the heads with the highest intervened logit difference in (a). We observe that only a small set of attention heads, and the first MLP layer exhibit strong causal influence on the model’s correct inference.
@@
Remark. In Figure 4 and the rest, unless otherwise specified, we adopt a calibrated version of the logit difference (see Appendix C): the closer to 1, the more significant the component is in “altering” the “belief” of the model on the selected subtask.
@@
Attention-head sub-component patching (QUERY-based patching). We now aim to understand why the attention heads identified in the last sub-section are important. For now, we continue with QUERY altering in the prompt pairs. Through intervening on the sub-components of each attention head, namely their value, key, and query, and through examining details of their attention weights, we find that there are roughly four types of attention heads. We show the results in Figure 5.
  1. nosep
@@
    Queried-rule locating head. Attention head (12,9)’s query activation has a large intervened logit difference according to Figure 5(a), therefore, its query and attention patterns are QUERY-dependent and contribute to altering the model’s “belief”. Furthermore, at the QUERY position, we find that on average, its attention weight is above 90% at the “conclusion” variable of the rule being queried. In other words, it is responsible for locating the queried rule, and storing that rule’s information at the QUERY position.555(9,25;26), (14,24;26) exhibit similar tendencies, albeit with smaller intervened logit differences.
  2. nosep
@@
    Queried-rule mover head. Attention head (13,11)’s value activations have large intervened logit difference, and intriguingly, its query and key activations do not share that tendency. This already suggests that its attention pattern performs a fixed action on both the original and altered prompts, and only the value information is sensitive to QUERY. Furthermore, within the relevant context (excluding the 6 in-context examples given), (13,11) assigns above 50% attention weight to the QUERY position, and its attention weight at QUERY is about 10 times larger than the second largest one on average. Recalling the role of layer 12, we find evidence that (13,11) moves the QUERY and queried-rule information to the “:” position.666(13,22), (15,8), (16,0) also appear to belong to this type, albeit with smaller intervened logit difference.
  3. nosep
@@
    Fact processing heads. Attention heads (16,12;14) and (14,26)’s query activations have large intervened logit differences. Within the relevant context, at the “:” token position, their attention weights are above 56, 80 and 63% respectively in the fact section of the context (starting from “Fact” and ending on “.” before “Question”).
  4. nosep
@@
    Decision head. Attention head (19,8)’s query activations have large intervened logit differences. Its attention pattern suggests that it is a “decision” head: within the relevant context, when the model is correct, the head’s top-2 attention weights are always on the correct starting node of the queried rule and the correct variable in the fact section, and the two token positions occupy more than 60% of its total attention in the relevant context on average. In other words, it already has the answer.777(17,25), (19,9;16) exhibit similar tendencies, albeit with smaller intervened logit differences.
@@
We delay detailed inspection and visualization of the attention statistics to Appendix C.3.3.
4.3.2 Circuit verification: a sufficiency test
A natural question now arises: is $\mathcal{C}$ sufficient to explain the (QUERY-sensitive) reasoning actions of the LLM? As explained in §4.2, we quantify sufficiency by measuring the “belief” altering effect of the circuit, and test $\mathcal{C}$ as follows. Given the original and altered prompt pairs $\{({\bm{X}}_{orig,n},{\bm{X}}_{alt,n})\}_{n\in[N]}$ which differ only by QUERY, we aim to verify that, simultaneously patching ${\bm{A}}_{\ell,h;t}({\bm{X}}_{orig,n})\to{\bm{A}}_{\ell,h;t}({\bm{X}}_{alt,n})$ for every attention head in the circuit $\mathcal{C}=\{(12,9),(13,11),(16,12),(19,8),...\}$ at their functioning token positions, while freezing all the other attention heads to the original activations ${\bm{A}}_{\ell,h;t}({\bm{X}}_{orig,n})$, lead to the average circuit-intervened logit difference $\Delta^{\mathcal{C}}_{orig\to alt}=\frac{1}{N}\sum_{n=1}^{N}\Delta_{orig\to alt,n}^{\mathcal{C}}$ approaching or surpassing the average logit difference of the (un-intervened) model run on the altered prompts, namely $\Delta_{alt}=\frac{1}{N}\sum_{n=1}^{N}\Delta_{alt,n}$. We confirm this hypothesis in Table 1 below.
$\mathcal{C}^{\dagger}$ $\mathcal{C}_{null}$ $\mathcal{C}$ $\mathcal{C}-QRLH$ $\mathcal{C}-QRMH$ $\mathcal{C}-FPH$ $\mathcal{C}-DH$
$\Delta^{\mathcal{C}^{\dagger}}_{orig\to alt}/\Delta_{alt}$1.0 0.98 -1.02 -0.99 -0.25 -0.89
Table 1: $\Delta^{\mathcal{C}^{\dagger}}_{orig\to alt}/\Delta_{alt}$, with different choices of $\mathcal{C}^{\dagger}$. $\mathcal{C}_{null}$ denotes the empty circuit, i.e. the case where no intervention is performed. We abbreviate the attention head families, for example, $DH=$ decision heads; $\mathcal{C}-DH=$ full circuit but with the decision heads removed.
To be precise, the circuit which we perform verification on is the union of the four attention head families, $\mathcal{C}=QRLH\cup QRMH\cup FPH\cup DH$ on their respective functioning token positions:
[…]
  • $DH=\text{Decision Heads}=\{(19,8;9;16),(17,25)\}$ patched at the “:” position.
We find that by patching all 14 attention heads in $\mathcal{C}$, $\Delta^{\mathcal{C}}_{orig\to alt}$ is about 98% of the “maximal” average logit difference $\Delta_{alt}$ on the altered samples. Moreover, removing any one of the four families of attention heads from $\mathcal{C}$ in the circuit interventions renders the “belief altering” effect of the intervention almost trivial. We present further discussions of the experimental procedure and results, and caveats of reasoning circuit verification in Appendix C.4.
4.3.3 Deeper examinations
In this subsection, we summarize the more involved experiments we performed in addition to the ones we presented so far, and refer the reader to their respective sections in the Appendix for further reading.
Early evidence of a similar reasoning circuit in Gemma-2-9B (§C.7). We performed a similar set of circuit discovery and verification experiments on Gemma-2-9B, and found that the set of attention heads in its circuit can also be divided into four families, closely resembling the attention patterns of the four families in Mistral-7B. In particular, Gemma-2-9B’s circuit is formed by the following families of attention heads, with $\mathcal{C}_{gemma}=QRLH\cup QRMH\cup FPH\cup DH$:
[…]
  • $DH=\text{Decision Heads}=\{(28,12),(30,9)\}$.
[…]
@@ 4.3.3 Deeper examinations
Further validating evidence of the attention head families in Mistral-7B (§C.5, C.6). We present further validating evidence of the queried-rule locating heads in §C.5, and the fact-processing and decision heads in §C.6, by performing causal intervention experiments in different sections of the input context.
  • For the queried-rule locating heads, we perform the following causal intervention experiments. In the original-altered prompt pairs, we only swap the location of the two rules in Rules section while keeping everything else the same. If these heads indeed perform their functions as we described, then when we run the model on the clean prompts, patching in the altered key activations at these heads (within the Rules section) should cause “negative” change to the model’s output, since it will cause these heads to mistake the queried-rule location in the altered prompt to be the right one, consequently storing the wrong rule information at the QUERY position. In particular, the model’s cross-entropy loss with respect to the original target should increase.
    Indeed, we observe in Figure 17 that interventions at these heads cause significantly larger increase in the cross-entropy loss than the other attention heads.
  • To further understand the fact-processing and decision heads, we construct original-altered pairs by only exchanging the truth value of the two nodes for the OR chain ([True, False]$\to$[False, True]). We perform subcomponent patching (query, key and value) of all attention heads in the Facts section, and found that out of all the patching results, only the fact-processing heads’ key activations exhibit strong “belief-altering” effects (large intervened logit difference). This indicates that, (1) the attention pattern of these heads are facts-sensitive, and (2) they perform (preliminary) filtering of the truth-value assignments, instead of blindly moving the truth-value assignments to the “:” position. These results are shown in Figure 18.
    Moreover, we also perform subcomponent patching at the “:” position, and found that only the decision heads’ query activations show nontrivial causal influence on the model’s correct output. This provides further evidence that these are the heads which relies on the embedding in the residual stream at the “:” position to send the correct answer tokens out. This result is shown in Figure 19.
@@ B.2.1 Transformer definition
The architecture definition follows that of GPT-2 closely. We illustrate the main components of this model in Figure 7, and point out where the frequently used terms in the main text of our paper are in this model.
Figure 7: Illustration of the major components of a 3-layer attention-only decoder-only transformer on the left, and a rough “sketch” of what is computed inside an attention block (2 attention heads for simplicity of the sketch).
@@
The following is the more technical definition of the model. Define input ${\bm{x}}=(x_{1},x_{2},...,x_{t})\in\mathbb{N}^{t}$, a sequence of tokens with length $t$. It is converted into a sequence of (trainable) token embeddings ${\bm{X}}_{token}=({\bm{e}}(x_{1}),{\bm{e}}(x_{2}),...,{\bm{e}}(x_{t}))^{T}\in\mathbb{R}^{t\times d_{e}}$, where we denote the hidden embedding dimension of the model with $d_{e}$. Adding to it the (trainable) positional embeddings ${\bm{P}}=({\bm{p}}_{1},{\bm{p}}_{2},...,{\bm{p}}_{t})^{T}\in\mathbb{R}^{t\times d_{e}}$, we form the zero-th layer embedding of the transformer
${\bm{X}}_{0}={\bm{X}}_{token}+{\bm{P}}=({\bm{e}}(x_{1})+{\bm{p}}_{1},...,{\bm{e}}(x_{t})+{\bm{p}}_{t}).$ (2)
This zero-th layer embedding is then processed by the attention blocks as follows.
Let the model have $L$ layers and $H$ heads. For layer index $\ell\in[L]$ and head index $j\in[H]$, attention head $\mathcal{A}_{\ell,j}$ is computed by
$\mathcal{A}_{\ell,j}({\bm{X}}_{\ell-1})=\mathcal{S}\left(\text{causal}\left[\frac{1}{\sqrt{d_{h}}}\left({\bm{Q}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}^{T}\right)^{T}{\bm{K}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}^{T}\right]\right)\widetilde{{\bm{X}}}_{\ell-1}{\bm{V}}_{\ell,j}^{T}\in\mathbb{R}^{t\times d_{h}},$ (3)
where $d_{h}=\frac{d_{e}}{H}$.
We explain how the individual components are computed below.
  • Let us begin with how the $\mathcal{S}(...)$ term is computed.
  • $\widetilde{{\bm{X}}}_{\ell-1}=\text{LayerNorm}({\bm{X}}_{\ell-1})\in\mathbb{R}^{t\times d_{e}}$, where LayerNorm denotes the layer normalization operator (Ba et al., 2016).
  • ${\bm{Q}}_{\ell,j},{\bm{K}}_{\ell,j}\in\mathbb{R}^{d_{h}\times d_{e}}$ are the key and query matrices of attention head $(\ell,j)$, where $d_{h}=\frac{d_{e}}{H}$. They are multiplied with the input $\widetilde{{\bm{X}}}_{\ell-1}$ to obtain the query and key activations ${\bm{Q}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}^{T}$ and ${\bm{K}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}$, both in the space $\mathbb{R}^{d_{h}\times t}$. We then perform the “scaled dot-product” of the query and key activations to obtain
    $\frac{1}{\sqrt{d_{h}}}\left({\bm{Q}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}^{T}\right)^{T}{\bm{K}}_{\ell,j},\widetilde{{\bm{X}}}_{\ell-1}^{T}\in\mathbb{R}^{t\times t},$ (4)
    which was introduced in Vaswani et al. (2017) and also used in GPT2 (Radford et al., 2019b).
  • The causal mask operator $\text{causal}:\mathbb{R}^{t\times t}\to\mathbb{R}^{t\times t}$ allows the lower triangular portion of the input (including the diagonal entries) to pass through unchanged, and sets the upper triangular portion of the input to $-U$, where $U$ is a very large positive number (some papers simply denote this $-U$ as $\infty$- ∞). In other words, given any ${\bm{M}}\in\mathbb{R}^{t\times t}$ and $(i,k)\in[t]\times[t]$,
    $\left[\text{causal}\left[{\bm{M}}\right]\right]_{i,k}=\left[{\bm{M}}\right]_{i,k},\,\text{if }i\geq k;$ (5)
    $\left[\text{causal}\left[{\bm{M}}\right]\right]_{i,k}=-U,\,\text{if }i<k.$
  • $\mathcal{S}:\mathbb{R}^{t\times t}\to\mathbb{R}^{t\times t}$ is the softmax operator, which computes the row-wise softmax output from the input square matrix. In particular, given a square input matrix ${\bm{M}}\in\mathbb{R}^{t\times t}$ with its upper triangular portion set to $-U$ (note that the causal mask operator indeed causes the input to $\mathcal{S}$ to have this property), we have
    $[\mathcal{S}({\bm{M}})]_{i,k}=\frac{\exp\left([{\bm{M}}]_{i,k}\right)}{\sum_{n=1}^{i}\exp([{\bm{M}}]_{i,n})},\,\text{if }i\geq k;$ (6)
    $[\mathcal{S}({\bm{M}})]_{i,k}=0,\,\text{if }i<k.$
  • To recap a bit, we have now explained how to compute the first major term in equation 3, namely $\mathcal{S}\left(\text{causal}\left[\frac{1}{\sqrt{d_{h}}}\left({\bm{Q}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}^{T}\right)^{T}{\bm{K}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}^{T}\right]\right)\in[0,1]^{t\times t}$. It reflects the attention pattern (also called attention probabilities) of the attention head $(\ell,j)$ illustrated in Figure 7’s right half. Intuitively speaking, the $(i,k)$ entry of this $t$ by $t$ matrix reflects how much the attention head moves the information from the previous layer $\ell-1$ at the source token position of $k$ to the current layer $\ell$ at the target token position $i$.
  • Now what about $\widetilde{{\bm{X}}}_{\ell-1}{\bm{V}}_{\ell,j}^{T}$? ${\bm{V}}_{\ell,j}\in\mathbb{R}^{d_{h}\times d_{e}}$ is the value matrix of attention head $(\ell,j)$. It is multiplied with $\widetilde{{\bm{X}}}_{\ell-1}$ to obtain the value activation $\widetilde{{\bm{X}}}_{\ell-1}{\bm{V}}_{\ell,j}^{T}\in\mathbb{R}^{t\times d_{h}}$.
  • At this point, we have shown how the whole term in equation equation 3 is computed.
Having computed the output of the all $H$ attention heads in the attention block at layer $\ell$, we find the output of the attention block as follows:
${\bm{X}}_{\ell}={\bm{X}}_{\ell-1}+\text{Concat}[\mathcal{A}_{\ell,1}({\bm{X}}_{\ell-1}),...,\mathcal{A}_{\ell,H}({\bm{X}}_{\ell-1})]{\bm{W}}_{O,\ell}^{T}.$ (7)
The operators are defined as follows:
  • $\text{Concat}[\cdot]$ is the concatenation operator, where $\text{Concat}[\mathcal{A}_{\ell,1}({\bm{X}}_{\ell-1}),...,\mathcal{A}_{\ell,H}({\bm{X}}_{\ell-1})]\in\mathbb{R}^{t\times d_{e}}$.
  • ${\bm{W}}_{O,\ell}\in\mathbb{R}^{d_{e}\times d_{e}}$ is the projection matrix (sometimes called output matrix) of layer $\ell$. In our implementation, we allow this layer to have trainable bias terms too.
Finally, having computed, layer by layer, the hidden outputs ${\bm{X}}_{\ell,t}$ for $\ell\in[L]$, we apply an affine classifier (with softmax) to obtain the output of the model
${\bm{f}}({\bm{x}})=\mathcal{S}(\widetilde{{\bm{X}}}_{L,t}{\bm{W}}_{class}^{T}+{\bm{b}}_{class})$ (8)
This output indicates the probability vector of the next word.
In this paper, we set the dimension of the hidden embeddings $d_{e}=768$.